# On the Preservation of Input/Output Directed Graph Informativeness under Crossover

### v1

@@ Abstract
There is a broad class of networks which connect inputs to outputs. While evolutionary operators have served as a backbone for a wide array of complex optimization problems, methods to apply such operators to these networks remain ill-defined. We aim to remedy this. We define an Input/Output Directed Graph (or IOD Graph) as a graph with a set of nodes $N$ and directed edges $E$, where $N$ contains (a) a set of “input nodes” $I\subset N$, where each $i\in I$ has no incoming edges and any number of outgoing edges, and (b) a set of “output nodes” $O\subset N$, where each $o\in O$ has no outgoing edges and any number of incoming edges, and $I\cap O=\emptyset$. We define informativeness, which involves the connections via directed paths from the input nodes to the output nodes: A partially informative IOD Graph has at least one path from an input to an output, a very informative IOD Graph has a path from every input to some output, and a fully informative IOD Graph has a path from every input to every output.
A perceptron is an example of an IOD Graph. If it has non-zero weights and any number of layers, it is fully informative. As links are removed (assigned zero weight), the perceptron might become very, partially, or not informative.
We define a crossover operation on IOD Graphs in which we find subgraphs with matching sets of forward and backward directed links to “swap.” With this operation, IOD Graphs can be subject to evolutionary computation methods. We show that fully informative parents may yield a non-informative child. We also show that under conditions of contiguousness and the no dangling nodes condition, crossover compatible, partially informative parents yield partially informative children, and very informative input parents with partially informative output parents yield very informative children. However, even under these conditions, full informativeness may not be retained.
@@ 1 Introduction
There is a broad class of networks which connect inputs to outputs. While evolutionary operators have served as a backbone for a wide array of complex optimization problems, methods to apply such operators to these networks remain ill-defined. We seek to remedy this.
We define this class of graphs as Input/Output Directed Graphs or IOD Graphs. An IOD Graph is a graph with a set of nodes $N$ and directed edges $E$, where $N$ contains (a) a set of “input nodes” $I\subset N$, where each $i\in I$ has no incoming edges and any number of outgoing edges, and (b) a set of “output nodes” $O\subset N$, where each $o\in O$ has no outgoing edges and any number of incoming edges, and $I\cap O=\emptyset$. Nodes $n\in N$ which are neither inputs nor outputs, so $n\notin I,n\notin O$, are called “intermediate nodes.”
IOD Graphs can represent a variety of flow or state transformation networks, any of which can be subject to crossover as we define it in this study. Examples of IOD Graphs include chemical reaction networks (e.g., Unsleber and Reiher, 2020; Wen, Spotte-Smith, Blau, McDermott, Krishnapriyan, and Persson, 2023), municipal water systems (e.g., Shinstine, Ahmed, and Lansey, 2002; Wu and Clark, 2009), data flow networks (e.g., Meng, Wong, Yuan, and Lu, 2004), and electrical circuits (e.g., Koza, Bennett, Andre, and Keane, 1997; Shanthi and Parthasarathi, 2009). A common use of IOD Graphs is problem solving. Multi-layer perceptrons, for example, are feed-forward IOD Graphs which convert inputs into “useful” outputs. Perceptrons are used to solve a variety of problems, for example in cognitive science they are used to emulate human categorization behavior. [^1] E.g. the classification learning literature beginning with Shepard, Hovland, and Jenkins (1961) and including the neural network ALCOVE Kruschke (1992). See more below. Because of the prominence of neural networks, we will generally describe IOD Graphs as information flow networks and interpret the properties and results from that perspective. However, as mentioned above, these results about crossover could also apply to e.g. municipal water system design. Also, while many neural networks are feed-forward, we do not restrict our study or the definition of IOD Graphs to feed-forward networks.
@@
One problem appropriate for IOD Graphs is the classification learning problem in cognitive science/psychology. A canonical version of this problem was by Shepard, Hovland, and Jenkins (1961). In their laboratory, they showed human subjects objects characterized by a three-dimensional binary vector (e.g. big v. small, dark v. light, square v. triangle) and queried the category, A or B. The experimenter knew the true category, and told the human subject whether their guess was right; the error rate over time of different categorizations was measured (as the reader might imagine, some categorizations are easier to learn than others) and these experiments provide a data benchmark that future computational learning models strived to explain. One such explanation was ventured by Kruschke (1992), who approached this problem with a single-layer perceptron/feed-forward neural network with three inputs, two outputs, and nine nodes in the “hidden layer,” for a total of $14$ nodes. The set of all IOD Graphs with $14$ nodes is vast: there are $18314^{2}-3^{2}-2^{2}=$ 183 possible connections, and therefore roughly $2^{183}\approx 1.2\times 10^{56}$ different possible IOD Graphs with that number of inputs, outputs, and intermediate nodes. Growing potential solutions in this space via evolutionary methods seems wise. The DIVA model (Kurtz, 2007) is another IOD Graph in this literature, which would model the SHJ problem as $3$ inputs and with two sets of $3$ outputs for a total of $6$ outputs. A crossover method like the one we describe in this paper could, for example, breed IOD Graphs from ALCOVE and DIVA implementations, possibly providing useful new solutions.
@@
The purpose of this paper is to rigorously define a crossover operation and (begin to) characterize how informativeness (and actionability) changes under this crossover operation. Crossover has been considered for feed-forward neural networks in some studies, e.g. Braun and Weisbrod (1993), Arifovic and Gencay (2001), García-Pedrajas, Ortiz-Boyer, and Hervás-Martínez (2006), Chandra, Frean, and Zhang (2012), Dragoni, Azzini, and Tettamanzi (2014), and Uriot and Izzo (2020). From that perspective, we seek to characterize instead of invent those methods. That is, we seek to lay a strong theoretical foundation for crossover on a broader class of graphs, of which feed-forward neural networks, municipal water systems, electric circuits, etc., are special cases, and relate these different methods to informativeness, which we believe helps capture some of the effectiveness of IOD Graphs to process information.
@@
We formally define IOD Graphs and informativeness in Section 2. We formally define the crossover operation in Section 3, including the core definitions (Section 3.1) and a discussion of crossover compatibility (Section 3.2), which characterizes when two IOD Graphs can be subject to the crossover operation. We then bring the analysis back to informativeness in Section 4, ‘The Preservation of Informativeness,’ which proves the core theorems. Section 5 briefly extends the analysis from informativeness to a related idea of actionability, which characterizes the amount of informed behavior flowing out of a system, in the sense that it characterizes how many outputs are eventually connected from an input. Section 6 discusses aspects of the results, such as the competing conventions problem and the distribution of informativeness categories across networks of different degree. Section 7 concludes.
@@ 2 Definitions
Figure 1(a) is a partially informative graph: For example, $I1\rightarrow V\rightarrow W\rightarrow Y\rightarrow Z\rightarrow O2$ is a directed path which leads from $I1\in I$ to $O2\in O$. However, it is not very informative, because there is no path from $I2$ to any output, and it is not fully informative because very informative is a necessary condition for fully informative. If the IOD Graph were modified by adding a directed edge from $I2\rightarrow W$, then this modified IOD Graph would be very informative. A multi-level perceptron with non-zero weights on its edges, for example depicted in Figure 1(b), is fully informative. Every input has a directed path to every output. Removing even one of these links, however, would reduce the IOD Graph depicted in Figure 1(b) to be only very informative.
@@ 3.1 Crossover-related Definitions
Given this definition of an IO partition, we define $M(\Psi{},\Omega{})$ as the set of all edges which connect the parts of the partition. This is called a cut in network theory. Formally, using $(n_{1},n_{2})$ to denote a directed edge from $n_{1}$ to $n_{2}$,
$M(\Psi{},\Omega{})$ =$=\left\{e\in E\quad\middle|\begin{array}[]{rl}e=&(n_{1},n_{2})\\ &\textrm{ or }\\ e=&(n_{2},n_{1})\end{array}\textrm{ for all }n_{1}\in\Psi{},n_{2}\in\Omega{}\right\}$
@@ 3.2 Crossover compatibility
Suppose $G,G^{\prime}$ are IOD Graphs with crossover compatible IO Partitions $(\Psi{},\Omega{})$ and $(\Psi{}^{\prime},\Omega{}^{\prime})$. Then a crossover membrane $\hat{M}$ is defined as as all edges $f^{\prime\prime}$ and $b^{\prime\prime}$, where
@@
Note that a given pair of crossover compatible IO Partitions $(\Psi{},\Omega{})$ and $(\Psi{}^{\prime},\Omega{}^{\prime})$, there must exist at least one crossover membrane. There may be, and often are, more than one possible crossover membrane. In that case, $\Psi{}$ and $\Omega{}^{\prime}$ can be connected in meaningfully different configurations.
@@ 4 The Preservation of Informativeness
It follows immediately that IOD Graphs which satisfy the No Dangling Nodes condition and have a non-empty set of intermediate nodes are partially informative. (Why? Any existing intermediate node must lie on a path from an input to an output, and therefore such a path exists.) However, an IOD Graph can satisfy the no dangling nodes condition and not be very informative, because there can be inputs which connect to nothing, as seen in Figure 6(b).
@@
In Figure 7, suppose that the node $n^{\prime\prime}$ is an arbitrary node in the input part $\Psi$. We seek to show that $n^{\prime\prime}$ must be on a path from an input to an output in the child $C$. In this example, that path turns out to be $i\rightarrow n^{\prime\prime}\rightarrow\sim n\rightarrow n^{\prime}\rightarrow o^{\prime}$. We build that path in pieces.
Consider the input parent $G$. Since $G$ is input-contiguous, we are assured of the path from input $i$ to node $n^{\prime\prime}$, which we call path $p$. Path $p$ lies entirely within $\Psi$. Since $G$ satisfies the No Dangling Nodes condition, we are also assured of a path from $i2$ to $O$. In this case, the path is $i2\rightarrow A\rightarrow n^{\prime\prime}\rightarrow\sim n\rightarrow B\rightarrow C\rightarrow O2$, which we call path $q$. This path was selected to illustrate that the path may cross between $\Psi$ and $\Omega$ multiple times; in particular, we are not assured from No Dangling Nodes alone that the subpath from $i2$ through $\sim n$ to $n^{\prime\prime}$ resides entirely within $\Psi$.
Now consider the output parent $G^{\prime}$ in Figure 7. Here, the No Dangling Nodes condition assures us that this $n^{\prime}\in\Omega{}$ that has been selected must have some path to some output $o^{\prime}$ that lies entirely within $\Omega$, which we call path $r$ [^8] Strictly speaking, No Dangling Nodes is not required for the Output Parent $G^{\prime}$ in this part of the proof. It is required later, when we consider $n^{\prime\prime}\in\Omega{}^{\prime}$.
@@
Now, because $G$ satisfies the no dangling nodes condition, $n^{\prime\prime}$ in $G$ must be on a path from $(\tilde{i},\ldots,n^{\prime\prime},\ldots,\tilde{o})$ for some $\tilde{i}\in I,\tilde{o}\in O$ [^9] Note that the initial segment of this path $(\tilde{i},\ldots,n^{\prime\prime})$ need not lie within $\Psi$, which is why the earlier path segment $p$ is needed. Consider the subpath $(n^{\prime\prime},\ldots,\tilde{o})$. Now that path must contain a node in $\Omega{}$. Find the first node in the path $(n^{\prime\prime},\ldots,\tilde{o})$ which is in $\Omega{}$, and consider the node $\tilde{n}$ which links to that node. We call this path $q=(i,\ldots,n^{\prime\prime},\ldots,\tilde{n})$. Note that $q$ resides entirely within $\Psi{}$. This means the path $p\rightarrow q$ resides entirely within $\Psi{}$.
@@
A symmetric argument applies, stated here for completeness: Since the IO Partition $(\Psi{}^{\prime},\Omega{}^{\prime})$ is contiguous, there must be a path from $n^{\prime\prime}$ to some $o^{\prime}\in O^{\prime}$ which is contained within $\Omega{}^{\prime}$. We shall call this path $p^{\prime}=(n^{\prime\prime},\ldots,o^{\prime})$. Now, $n^{\prime\prime}$ in $G^{\prime}$ is on a path from $(\bar{i},\ldots,n^{\prime\prime},\ldots,\bar{o})$ for some $\bar{i}\in I^{\prime},\bar{o}\in O^{\prime}$. Consider the path $(\bar{i},\ldots,n^{\prime\prime})$. Now that path must contain a node in $\Psi{}^{\prime}$. Find the last node in the path $(\bar{i},\ldots,n^{\prime\prime})$ which is in $\Psi{}$, and consider the node $\bar{n}$ which links from that node. We name this path $q^{\prime}=(\bar{n},\ldots,n^{\prime\prime},\ldots,\bar{o})$. Note that $q$ resides entirely within $\Omega{}^{\prime}$ and therefore $q^{\prime}\rightarrow p^{\prime}$ also resides entirely within $\Omega{}$. In $C$, there must be some some $n\in\Psi{}$ which links to $\bar{n}$. By the IO partition $(\Psi{},\Omega{})$ being contiguous, there must be a path an input $i\in I$ to $n$ fully contained in $\Psi{}$, which we call $r^{\prime}=(i,\ldots,n)$. Therefore the path

### v3

@@ Abstract
There exists a broad class of networks that connect inputs to outputs. These networks include chemical transformation networks, electrical circuits, municipal water systems, and neural networks. The goals of this paper are to provide a theoretical foundation for evolutionary crossover on this class of graphs and connect crossover to informativeness, a measure of the connectedness of inputs to outputs. Informativeness is defined as: a partially informative graph has at least one path from an input to some output, a very informative graph has a path from every input to some output, and a fully informative graph has a path from every input to every output. If a neural network with non-zero weights and any number of layers is fully informative. As links are removed (assigned zero weight), it may become very, partially, or not informative. (The complement of informativeness is actionability, which is a measure of how connected outputs are from inputs.)
We define a crossover operation on IOD Graphs in which we find subgraphs with matching sets of forward and backward directed links to “swap.” With this operation, IOD Graphs can be subject to evolutionary computation methods. We show that fully informative parents may yield a non-informative child. We also show that under certain conditions, crossover compatible, partially informative parents yield partially informative children, and very informative input parents with partially informative output parents yield very informative children. However, even under these conditions, full informativeness may not be retained. Similar results hold for actionability.
@@ 1 Introduction
There is a broad class of networks which connect inputs to outputs, used for modeling and problem solving in a variety of domains. Examples include chemical reaction networks (e.g., Unsleber and Reiher,, 2020; Wen et al.,, 2023), municipal water systems (e.g., Shinstine et al.,, 2002; Wu and Clark,, 2009), data flow networks (e.g., Meng et al.,, 2004), and electrical circuits (e.g., Koza et al.,, 1997; Shanthi and Parthasarathi,, 2009). We call these graphs Input/output Directed Graphs or IOD Graphs. A common IOD Graph is the multi-layer perceptron, which is a feed-forward IOD Graph that converts inputs into “useful” outputs. Perceptrons are used to solve a variety of problems; for example, in cognitive science they are used to emulate human categorization behavior. [^1] E.g., the classification learning literature beginning with Shepard et al., (1961) and including the neural network ALCOVE (Kruschke,, 1992). See more below.
While evolutionary operations like crossover have been used in all these domains, the application of crossover to IOD Graphs remains sparse. We provide a theoretical foundation for crossover across this class of networks. First, we establish how crossover can be applied to any two ore more such networks. We also connect crossover to informativeness, a measure of the connectedness of inputs to outputs, and provide proofs putting bounds on how informativeness flows from parent networks to their children.
First, we formally define this class of graphs as Input/Output Directed Graphs or IOD Graphs: An IOD Graph is a graph with a set of nodes $N$ and directed edges $E$, where $N$ contains (a) a set of “input nodes” $I\subset N$, where each $i\in I$ has no incoming edges and any number of outgoing edges, and (b) a set of “output nodes” $O\subset N$, where each $o\in O$ has no outgoing edges and any number of incoming edges, and $I\cap O=\emptyset$. Nodes $n\in N$ which are neither inputs nor outputs, so $n\notin I,n\notin O$, are called “intermediate nodes.”
Because of the prominence of neural networks, we will generally describe IOD Graphs as information flow networks and interpret the properties and results from that perspective. However, as mentioned above, these results about crossover could also apply to e.g. municipal water system design. Also, while many neural networks are feed-forward, meaning the networks do not contain any feedback loops, we do not restrict our study or the definition of IOD Graphs to feed-forward networks.
@@
One problem appropriate for IOD Graphs is the classification learning problem in cognitive science/psychology. A canonical version of this problem was by Shepard et al., (1961). In their laboratory, they showed human subjects objects characterized by a three-dimensional binary vector (e.g. big v. small, dark v. light, square v. triangle) and queried the category, A or B. The experimenter knew the true category, and told the human subject whether their guess was right; the error rate over time of different categorizations was measured (as the reader might imagine, some categorizations are easier to learn than others) and these experiments provide a data benchmark that future computational learning models strived to explain. One such explanation was ventured by Kruschke, (1992), who approached this problem with a single-layer perceptron/feed-forward neural network with three inputs, two outputs, and nine nodes in the “hidden layer,” for a total of $14$ nodes. The set of all IOD Graphs with $14$ nodes is vast: assuming at most one link between nodes, there are roughly $2^{132}\approx 5.4\times 10^{39}$ different possible IOD Graphs with that number of inputs, outputs, and intermediate nodes. Growing potential solutions in this space via evolutionary methods seems wise. The DIVA model (Kurtz,, 2007) is another IOD Graph in this literature, which would model the SHJ problem as $3$ inputs and with two sets of $3$ outputs for a total of $6$ outputs. A crossover method like the one we describe in this paper could, for example, breed IOD Graphs from ALCOVE and DIVA implementations, possibly providing useful new solutions.
@@
The purpose of this paper is to rigorously define a crossover operation and (begin to) characterize how informativeness and actionability changes under this crossover operation. That is, we seek to lay a strong theoretical foundation for crossover on a broader class of graphs, of which feed-forward neural networks, municipal water systems, electric circuits, etc., are special cases, and relate these different methods to informativeness, which we believe helps capture some of the effectiveness of IOD Graphs to process information.
Since all neural networks are IOD Graphs, this paper is contribution to the substantial research effort applying crossover to neural networks. Starting with the perceptron (Rosenblatt,, 1957), the vast majority of neural networks explored, including modern deep nets, are feed-forward networks. While more sophisticated computation is possible when networks include recurrent links, there has yet to emerge a comprehensive theory of how the computational power of networks with recurrent links can be achieved. There has been much empirical research aimed at evolutionary computation for this task; for a recent summary, see Stanley et al., (2019).
The crossover process described here is applied on the graph directly and does not use evo-devo. Evo-devo (or Evolutionary Development) is a gene-driven process in which a genome representation is translated into phenotypes, in this case, graphs, and crossover is done on the genomes, not the graphs themselves (Jacob,, 1977; Gould,, 1977). This direct crossover on network structure may foster Lamarckian evolution and may be compatible with some evo-devo crossover methods (see Section 6.1).
@@
We formally define IOD Graphs and informativeness in Section 2. We formally define the crossover operation in Section 3, including the core definitions (Section 3.1) and a discussion of crossover compatibility (Section 3.2), which characterizes when two IOD Graphs can be subject to the crossover operation. We apply this crossover operation to informativeness in Section 4, ‘The Preservation of Informativeness,’ which proves the core theorems. Section 5 briefly extends the analysis from informativeness to a related idea of actionability, which characterizes the amount of informed behavior flowing out of a system, in the sense that it characterizes how many outputs are connected from an input. Section 6 discusses aspects of the results, such as the competing conventions problem and the distribution of informativeness categories across networks of different degree. Section 7 concludes.
@@ 2 Definitions
Figure 1(a) is a partially informative graph: For example, $I1\rightarrow V\rightarrow W\rightarrow Y\rightarrow Z\rightarrow O2$ is a directed path which leads from $I1\in I$ to $O2\in O$. However, it is not very informative, because there is no path from $I2$ to any output, and it is not fully informative because very informative is a necessary condition for fully informative. If the IOD Graph were modified by adding a directed edge from $I2\rightarrow W$, then this modified IOD Graph would be very informative. A multi-level perceptron with non-zero weights on its edges, for example depicted in Figure 1(b), is fully informative. Every input has a directed path to every output.
@@ 3.1 Crossover-related Definitions
Given this definition of an IO partition, we define $M(\Psi{},\Omega{})$ as the set of all edges which connect the parts of the partition. This is called a cut in network theory. Using $(n_{1},n_{2})$ to denote a directed edge from $n_{1}$ to $n_{2}$ and E to denote the list of all edges, we can formally define $M(\Psi{},\Omega{})$ as follows:
For all $e\in E$, where $e=(n_{1},n_{2})$ or $e=(n_{2},n_{1})$
$M(\Psi{},\Omega{})$ =$=\left\{e\quad\middle|\quad n_{1}\in\Psi,n_{2}\in\Omega\right\}$
@@ 3.2 Crossover compatibility
Suppose $G,G^{\prime}$ are IOD Graphs with crossover compatible IO Partitions $(\Psi{},\Omega{})$ and $(\Psi{}^{\prime},\Omega{}^{\prime})$. Then a $(G,G^{\prime})$ crossover membrane $\hat{M}$ is defined as as all edges $f^{\prime\prime}$ and $b^{\prime\prime}$, where
@@
Note that a given pair of crossover compatible IO Partitions $(\Psi{},\Omega{})$ and $(\Psi{}^{\prime},\Omega{}^{\prime})$, there must exist at least one crossover membrane, as shown in Lemma 1. There may be, and often are, more than one possible crossover membrane. In that case, $\Psi{}$ and $\Omega{}^{\prime}$ can be connected in meaningfully different configurations.
Lemma 1
Suppose $G,G^{\prime}$ are two crossover compatible IOD Graphs. Then there exists a $(G,G^{\prime})$ crossover membrane.
Proof. Since $G,G^{\prime}$ are crossover compatible, then select crossover compatible IO Partitions $(\Psi{},\Omega{})$ for $G$ and $(\Psi{}^{\prime},\Omega{}^{\prime})$ for $G^{\prime}$.
Then define $f,f^{\prime}$ and $b,b^{\prime}$ as the sets of forward and backward links; namely let $f,f^{\prime}$ be $f=F(\Psi{},\Omega{})$ and $f^{\prime}=F(\Psi{}^{\prime},\Omega{}^{\prime})$; and let $b,b^{\prime}$ be $b=B(\Psi{},\Omega{})$ and $b^{\prime}=B(\Psi{}^{\prime},\Omega{}^{\prime})$.
By definition of crossover compatible, $|f|=|f^{\prime}|$ and $|b|=|b^{\prime}|$. Create two one-to-one and onto mappings $g:f\rightarrow f^{\prime}$ and $h:b^{\prime}\rightarrow b$.
For a link $l$, define ${l}_{\textrm{source}}$ as the source node of $l$ and let ${l}_{\textrm{dest}}$ as the destination node of $l$; so by definition each link $l=({1}_{\textrm{source}},{l}_{\textrm{dest}})$.
Then define $\hat{M}_{f}$, $\hat{M}_{b}$ as:
$\hat{M}_{f}$ =$=\left\{({x}_{\textrm{source}},{g(x)}_{\textrm{dest}})\quad\middle|\quad x\in f\right\}$
$\hat{M}_{b}$ =$=\left\{({y}_{\textrm{source}},{h(y)}_{\textrm{dest}})\quad\middle|\quad y\in b^{\prime}\right\}$
$\hat{M}_{f}\cup\hat{M}_{b}$ is a $(G,G^{\prime})$ crossover membrane.
Now we define the child of crossover:
@@
With this definition in hand, we can define IOD Graph crossover
Definition 7
IOD Graph Crossover is any algorithm which transforms any two IOD Graphs with crossover compatible IO Partitions into a crossover child.
@@ 4 The Preservation of Informativeness
It follows immediately that IOD Graphs which satisfy the No Dangling Nodes condition and have a non-empty set of intermediate nodes are partially informative. This is because any existing intermediate node must lie on a path from an input to an output, and therefore such a path exists. However, an IOD Graph can satisfy the no dangling nodes condition and not be very informative, because there can be inputs which connect to nothing, as seen in Figure 6(b).
@@
In Figure 7, suppose that the node $n^{\prime\prime}$ is an arbitrary node in the input part $\Psi$. We seek to show that $n^{\prime\prime}$ must be on a path from an input to an output in the child $C$. In this example, that path turns out to be $i\rightarrow n^{\prime\prime}\rightarrow\sim n\rightarrow n^{\prime}\rightarrow o^{\prime}$. We build that path in pieces: pieces $p$, $q$, and $r$, which we define below.
Consider the input parent $G$. Since $G$ is input-contiguous, we are assured of the path from input $i$ to node $n^{\prime\prime}$, which we call path $p$. Path $p$ lies entirely within $\Psi$. Since $G$ satisfies the No Dangling Nodes condition, we are also assured of a path from $i2$ to $O$. In this case, the path is $i2\rightarrow A\rightarrow n^{\prime\prime}\rightarrow\sim n\rightarrow B\rightarrow C\rightarrow O2$. This path was selected to illustrate that the path may cross between $\Psi$ and $\Omega$ multiple times; in particular, we are not assured from No Dangling Nodes alone that the subpath from $i2$ through $\sim n$ to $n^{\prime\prime}$ resides entirely within $\Psi$.
There is a part of the path $i2\rightarrow A\rightarrow n^{\prime\prime}\rightarrow\sim n\rightarrow B\rightarrow C\rightarrow O2$ which we use later: the subpath $n^{\prime\prime}\rightarrow\sim n$, which we call path $q$.
Now consider the output parent $G^{\prime}$ in Figure 7. Here, the No Dangling Nodes condition assures us that this $n^{\prime}\in\Omega{}^{\prime}$ that has been selected must have some path to some output $o^{\prime}$ that lies entirely within $\Omega$’, which we call path $r$ [^8] Strictly speaking, No Dangling Nodes is not required for the Output Parent $G^{\prime}$ in this part of the proof. It is required later, when we consider $n^{\prime\prime}\in\Omega{}^{\prime}$.
@@
Now, because $G$ satisfies the no dangling nodes condition, $n^{\prime\prime}$ in $G$ must be on a path from $(\tilde{i},\ldots,n^{\prime\prime},\ldots,\tilde{o})$ for some $\tilde{i}\in I,\tilde{o}\in O$ [^9] Note that the initial segment of this path $(\tilde{i},\ldots,n^{\prime\prime})$ need not lie within $\Psi$, which is why the earlier path segment $p$ is needed. Consider the subpath $(n^{\prime\prime},\ldots,\tilde{o})$. Now that path must contain a node in $\Omega{}$. Find the first node in the path $(n^{\prime\prime},\ldots,\tilde{o})$ which is in $\Omega{}$, and consider the node $\tilde{n}$ which links to that node. That is the path $(i,\ldots,n^{\prime\prime},\ldots,\tilde{n})$. We define the subpath $q=(n^{\prime\prime},\ldots,\tilde{n})$. Note that $q$ resides entirely within $\Psi{}$. This means the path $p\rightarrow q$ resides entirely within $\Psi{}$.
@@
A symmetric argument applies, stated here for completeness: Since the IO Partition $(\Psi{}^{\prime},\Omega{}^{\prime})$ is contiguous, there must be a path from $n^{\prime\prime}$ to some $o^{\prime}\in O^{\prime}$ which is contained within $\Omega{}^{\prime}$. We shall call this path $p^{\prime}=(n^{\prime\prime},\ldots,o^{\prime})$. Now, $n^{\prime\prime}$ in $G^{\prime}$ is on a path from $(\bar{i},\ldots,n^{\prime\prime},\ldots,\bar{o})$ for some $\bar{i}\in I^{\prime},\bar{o}\in O^{\prime}$. Consider the path $(\bar{i},\ldots,n^{\prime\prime})$. Now that path must contain a node in $\Psi{}^{\prime}$. Find the last node in the path $(\bar{i},\ldots,n^{\prime\prime})$ which is in $\Psi{}$, and consider the node $\bar{n}$ which links from that node. This is the path $(\bar{n},\ldots,n^{\prime\prime},\ldots,\bar{o})$. We name the subpath $q^{\prime}=(\bar{n},\ldots,n^{\prime\prime})$. Note that $q^{\prime}$ resides entirely within $\Omega{}^{\prime}$ and therefore $q^{\prime}\rightarrow p^{\prime}$ also resides entirely within $\Omega{}$. In $C$, there must be some some $n\in\Psi{}$ which links to $\bar{n}$. By the IO partition $(\Psi{},\Omega{})$ being contiguous, there must be a path an input $i\in I$ to $n$ fully contained in $\Psi{}$, which we call $r^{\prime}=(i,\ldots,n)$. Therefore the path